OpenStudy (uscrnamc):

What is the point slope form of the line with slope -3/7 that passes through the point (5,8) THE ANSWER IS C

OpenStudy (uscrnamc):

A. y + 5 = -3/7 (x + 8) B. y + 8 = -3/7 (x +5) C. y - 8 = -3/7 (x - 5) D. y - 5 = -3/7 (x - 8)

OpenStudy (uscrnamc):

FOR ANYONE WANTING THE ANSWER FOR LATER ON IT IS C